OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cos(5pi/12)

OpenStudy (tanya123):

os 5π/12 cos ((5 pi/6)/2) = +/- sqrt( (1+cos (5 pi/6))/2 ) cos(5 pi/6) = -sqrt(3)/2 cos (5 pi/12) = +/- sqrt( (2/2-sqrt(3)/2)/2 ) cos (5 pi/12) = +/- sqrt( ((2-sqrt(3))/2)/2 ) cos (5 pi/12) = sqrt( ( (2-sqrt(3))/4 ) ) 5 pi/12 is in quadrant 1 so we choose the + sign because cos is positive there

OpenStudy (jdoe0001):

\(\bf \cfrac{5}{12}\to \cfrac{\cancel{ 3 }}{\cancel{ 12 }}+\cfrac{\cancel{ 2 }}{\cancel{ 12 }}\to \cfrac{1}{4}+\cfrac{1}{6}\qquad thus \\ \quad \\ \cfrac{5\pi}{12}\to \cfrac{1\pi}{4}+\cfrac{1\pi}{6}\to \cfrac{\pi}{4}+\cfrac{\pi}{6}\qquad thus \\ \quad \\ cos\left(\frac{5\pi}{12}\right)\to cos\left(\frac{\pi}{4}+\frac{\pi}{6}\right)\)

OpenStudy (jdoe0001):

hhh ohh the half angle... shoot... so... check above then

OpenStudy (anonymous):

ok, thank you!!